OpenStudy (anonymous):

write sin x + sin 5x as a product

OpenStudy (anonymous):

= 2 sin 'semi-sum' cos 'semi-difference' = 2 sin 3x cos 2x

OpenStudy (anonymous):

\( \sin A + \sin B = 2 \sin \frac{A+B}{2} \cos \frac {A-B}{2}, where , A>B \)

OpenStudy (anonymous):

This is one of the Prosthaphaeresis Formulas ( http://mathworld.wolfram.com/ProsthaphaeresisFormulas.html)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Glad to help :)